# Conformal Prediction in Dynamic Biological Systems

### v1

@@ Abstract
Uncertainty quantification (UQ) is the process of systematically determining and characterizing the degree of confidence in computational model predictions. In the context of systems biology, especially with dynamic models, UQ is crucial because it addresses the challenges posed by nonlinearity and parameter sensitivity, allowing us to properly understand and extrapolate the behavior of complex biological systems. Here, we focus on dynamic models represented by deterministic nonlinear ordinary differential equations. Many current UQ approaches in this field rely on Bayesian statistical methods. While powerful, these methods often require strong prior specifications and make parametric assumptions that may not always hold in biological systems. Additionally, these methods face challenges in domains where sample sizes are limited, and statistical inference becomes constrained, with computational speed being a bottleneck in large models of biological systems. As an alternative, we propose the use of conformal inference methods, introducing two novel algorithms that, in some instances, offer non-asymptotic guarantees, enhancing robustness and scalability across various applications. We demonstrate the efficacy of our proposed algorithms through several scenarios, highlighting their advantages over traditional Bayesian approaches. The proposed methods show promising results for diverse biological data structures and scenarios, offering a general framework to quantify uncertainty for dynamic models of biological systems.The software for the methodology and the reproduction of the results is available at https://zenodo.org/doi/10.5281/zenodo.13644870.
@@ 3.4 Case IV: NFKB signaling pathway
The dataset generation process was similar to the one used in the previous case studies. The results of applying different algorithms to a 13-point synthetic dataset are illustrated in the Supplementary Information. As shown, both methods based on conformal inference yield results that are in close agreement with each other. However, in this case, the computational cost difference becomes particularly significant. While the CUQDyn1 and CUQDyn2 algorithms compute the regions within a few minutes, the Bayesian approach requieres several hours to achieve the same task.
@@ Acknowledgments
JRB acknowledges support from grant PID2020-117271RB-C22 (BIODYNAMICS) funded by MCIN/AEI/10.13039/01100011033, and from CSIC grant PIE 202470E108 (LARGO). The authors also wish to thank Javier Enrique Aguilar Romero for his assistance with the use of Stan. The funders had no role in the study design, data collection and analysis, decision to publish, or preparation of the manuscript
@@ Appendix A Matlab implementation of the algorithms
The software for the methodology and the reproduction of the results is available at https://zenodo.org/doi/10.5281/zenodo.13644870. All computations were carried out on a PC DELL Precision 7920 workstation with dual Intel Xeon Silver 4210R processors.
@@ C.4 Case IV: NFKB signaling pathway
As shown, both methods based on conformal inference yield results that are in close agreement with each other. However, in this case, the computational benefits of our conformal strategies are especially pronounced. The CUQDyn1 and CUQDyn2 algorithms can compute the regions in just a few minutes, whereas the Bayesian approach requires several hours to accomplish the same task.

### v2

@@ Abstract
Uncertainty quantification (UQ) is the process of systematically determining and characterizing the degree of confidence in computational model predictions. In the context of systems biology, especially with dynamic models, UQ is crucial because it addresses the challenges posed by nonlinearity and parameter sensitivity, allowing us to properly understand and extrapolate the behavior of complex biological systems. Here, we focus on dynamic models represented by deterministic nonlinear ordinary differential equations. Many current UQ approaches in this field rely on Bayesian statistical methods. While powerful, these methods often require strong prior specifications and make parametric assumptions that may not always hold in biological systems. Additionally, these methods face challenges in domains where sample sizes are limited, and statistical inference becomes constrained, with computational speed being a bottleneck in large models of biological systems. As an alternative, we propose the use of conformal inference methods, introducing two novel algorithms that, in some instances, offer non-asymptotic guarantees, enhancing robustness and scalability across various applications. We demonstrate the efficacy of our proposed algorithms through several scenarios, highlighting their advantages over traditional Bayesian approaches. The proposed methods show promising results for diverse biological data structures and scenarios, offering a general framework to quantify uncertainty for dynamic models of biological systems.The software for the methodology and the reproduction of the results is available at https://zenodo.org/doi/10.5281/zenodo.13644869.
@@ 3.4 Case IV: NFKB signaling pathway
The dataset generation process was similar to the one used in the previous case studies. The results of applying different algorithms to a 13-point synthetic dataset are illustrated in the Supplementary Information. Both conformal inference methods yielded accurate and consistent results in short computation times. In contrast, STAN failed to converge, even after very long computations. This non-convergence is likely due to partial lack of identifiability.
@@ Acknowledgments
JRB acknowledges support from grant PID2020-117271RB-C22 (BIODYNAMICS) funded by MCIN/AEI/10.13039/501100011033, from grant PID2023-146275NB-C22 (DYNAMO-bio) funded by MICIU/AEI/ 10.13039/501100011033 and ERDF/EU, and from grant CSIC PIE 202470E108 (LARGO). The authors also wish to thank Javier Enrique Aguilar Romero for his assistance with the use of Stan. The funders had no role in the study design, data collection and analysis, decision to publish, or preparation of the manuscript
@@ Appendix A Matlab implementation of the algorithms
The software for the methodology and the reproduction of the results is available at https://zenodo.org/doi/10.5281/zenodo.13644869. All computations were carried out on a PC DELL Precision 7920 workstation with dual Intel Xeon Silver 4210R processors.
@@ C.4 Case IV: NFKB signaling pathway
Figure 8: Comparative analysis of the NFkb signaling pathway model predictive regions. This figure presents the $95\%$ predictive regions obtained from a 13-point synthetic dataset. It showcases the regions for the six observables obtained by using our two proposed methods: CUQDyn1 and CUQDyn2.
@@
Figure 8 shows the results of applying our two methods to a 13-point synthetic dataset. Both methods based on conformal inference yielded results that are in close agreement with each other. However, in this case, we were not able to obtain adequate predictive regions using STAN, even after many hours of computation, probably due to the partial lack of identifiability. Remarkably, our CUQDyn1 and CUQDyn2 algorithms can compute the regions in just a few minutes using a standard PC.
The dataset generation process was the same that the one used in the previous case studies.